# Quaternion tensor low rank approximation

### v2

@@ 6 Experiments
$PSNR:=10\log_{10}\left(\frac{H\times W\times T\times||\dot{\mathcal{X}}||_{\infty}^{2}}{||\dot{\mathcal{Q}}-\dot{\mathcal{X}}||_{F}^{2}}\right),$
@@
The convergence criteria for Algorithm 4 includes conditions on the Frobenius norm of the differences between successive iterations: $||\dot{\mathcal{Q}}^{(\tau+1)}-\dot{\mathcal{Q}}^{(\tau)}||_{F}^{2}<tol$, $||\dot{\mathcal{S}}^{(\tau+1)}-\dot{\mathcal{S}}^{(\tau)}||_{F}^{2}<tol$, and $||\dot{\mathcal{P}}^{(\tau+1)}-\dot{\mathcal{P}}^{(\tau)}||_{F}^{2}<tol$ for Algorithm 3, where the tolerance $tol$ is set to $1\times 10^{-6}$.
@@ 6.1 Completion task
To demonstrate the effectiveness of the methods, we use the sample rate (SR), which indicates the percentage of missing pixels, chosen from multiple levels: ${0.1,0.3,0.5}$. A higher sample rate corresponds to a greater number of omitted pixels and, consequently, a more challenging task for the methods. The index set for the missing values is the same across the three channels but differs for each frame, which increases the difficulty of the problem.
@@ 6.2 Denoising task
[…]

### v3

@@ 6 Experiments
[…]
@@
The convergence criteria for Algorithm 4 includes conditions on the Frobenius norm of the differences between successive iterations: $\dot{\mathcal{Q}}^{(\tau+1)}-\dot{\mathcal{Q}}^{(\tau)}||_{F}^{2}<tol$, $\dot{\mathcal{S}}^{(\tau+1)}-\dot{\mathcal{S}}^{(\tau)}||_{F}^{2}<tol$, and $\dot{\mathcal{P}}^{(\tau+1)}-\dot{\mathcal{P}}^{(\tau)}||_{F}^{2}<tol$ for Algorithm 3, where the tolerance toltol is set to $1\times 10^{-6}$.
@@ 6.1 Completion task
To demonstrate the effectiveness of the methods, we use the sample rate (SR), which indicates the percentage of missing pixels, chosen from multiple levels: 0.1,0.3,0.5{0.1,0.3,0.5}. A higher sample rate corresponds to a greater number of omitted pixels and, consequently, a more challenging task for the methods. The index set for the missing values is the same across the three channels but differs for each frame, which increases the difficulty of the problem.
@@ 6.2 Denoising task
Figure 3: Visual results of different methods on Noise level $\%$, on a frame image from Highway Dataset. The forth and fifth column, are TRPCA-NC-rand, and TRPCA-NC-dct, respectively.